OpenStudy (anonymous):

Find the area of the triangle with the given measurements. Round the solution to the nearest hundredth if necessary. B = 74°, a = 14 cm, c = 20 cm

OpenStudy (anonymous):

c

OpenStudy (jdoe0001):

have you covered the Law of Cosines yet?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

ok... so we would want to find the side "b" then so \(\bf \textit{Law of Cosines}\\ \quad \\ b^2 = {\color{blue}{ a}}^2+{\color{red}{ c}}^2-(2{\color{blue}{ a}}{\color{red}{ c}})cos(B)\implies b = \sqrt{{\color{blue}{ a}}^2+{\color{red}{ c}}^2-(2{\color{blue}{ a}}{\color{red}{ c}})cos(B)} \\ \quad \\ thus\implies b=\sqrt{14^2+20^2-(2\cdot 14\cdot 20)\cdot cos(74^o)}\)

OpenStudy (anonymous):

b=21.0153

OpenStudy (jdoe0001):

yeap.... so now we know the sides "a", "b" and "c" so to get the area of a triangle, with only the sides length we could just use Heron's Formula, that is \(\bf {\color{purple}{ s}}=\cfrac{a+b+c}{2} \\ \quad \\ \textit{area of a triangle}=\sqrt{{\color{purple}{ s}}({\color{purple}{ s}}-a)({\color{purple}{ s}}-b)({\color{purple}{ s}}-c)}\)

OpenStudy (anonymous):

Thanks, can you help me check something else?

OpenStudy (jdoe0001):

I'd be dashing soon... so is better to post anew anyhow... so if I dunno... someone else may

OpenStudy (anonymous):

ok